OpenStudy (anonymous):

how the work to solve |x+3| +7 ≤ 10

OpenStudy (anonymous):

Also, Describe the graph of the solution in a complete sentence. Please (:

OpenStudy (anonymous):

Let's first isolate the absolute value term. \[|x+3| \le 3\]To eliminate the absolute value, we need to write two equations. One for the positive case, and another for the negative case. \[x+3 \le3 ~~\text{and}~~ x+3 \ge -3\]It can be seen that this simplifies to \[-3 \le x+3 \le 3\]Subtracting through by 3, we get \[-6 \le x \le 0\]A plot can be seen here: http://www.wolframalpha.com/input/?i=abs%28x%2B3%29+%2B+7+%3C%3D+10

OpenStudy (anonymous):

eashmore thanxx for graph .........i got its website which is very useful

OpenStudy (anonymous):

@eashmore ..do u hav any other website

OpenStudy (anonymous):

You're welcome! Indeed, Wolfram Alpha is a wonderful site.

OpenStudy (anonymous):

http://mathworld.wolfram.com is great. It discusses all kinda of math concepts. A great reference. I find it easier to read than wikipedia.

OpenStudy (anonymous):

thnxxxx i m reading thos web sites its wonderful......onece again thnxxxx

OpenStudy (anonymous):

thanx eashmore for the useful websites.. do u know know any book or site of calculus??

OpenStudy (anonymous):

Wolfram|Alpha can do calculus calculations, but it doesn't show how it computs answers very well. If you do a Google search of what you're trying to work through, you'll find that a lot of notes are out there from various universities. I often find these helpful, and are a great addition to the class notes you already have. http://calculus.org it a pretty good compilation of calculus resources.

OpenStudy (anonymous):

thanx.. i visited mathworld.wolfram... its amazing thanx again:)